The firm has net income of $3450 and total assets of $13,800. The payout ratio is 30 percent. What is thoe internal growth rate? a. 14.5% b. 17.896 21.3% 29,4% e. 33.3% C.
Answer #1

Return on Assets =Net Income/Total Assets =3450/13800=25%

Internal growth Rate=(1-Dividend Payout)*Return on Assets

=(1-30%)*(25%)

=17.5%